# The ubiquity of variable radio emission and spin-down rates in pulsars

### v2

@@ 4.2 New pulsars with both profile and spin-down variations
The radio profile of this pulsar is has two peaks that we found to exhibit substantial variability. The leading component shows both fast and slow changes, while the trailing component only varies slowly with time. In contrast, the spin-down rate varies smoothly over time, with a potential quasi-period of 1648 d. Our correlation map suggests the profile variations lag the changing spin-down rate by 500-1000 d.
@@ 5.1 Scaling relations across the population
While a direct causal relationship between the observed spin-down variability in pulsars and external or internal physical processes is yet to be established, we can explore how this behaviour scales phenomenologically with various pulsar properties. In Figure 12 we present the measured min/max difference in spin-down rate against the corresponding ‘weak’ (minimum) spin-down rate value for the 238 variable pulsars in our sample, alongside values from several previous works. Lyne et al. (2010) noted their sample of 17 correlated profile/spin-down state switching pulsars appeared to follow a simple power-law relation, where . A somewhat shallower power-law relation of was recovered using extended timing of the same pulsar sample by Shaw et al. (2022). While there is a substantial number of pulsars in Figure 12 that reside about the original Lyne et al. (2010) relation (shown in black), about half of our larger sample are situated between one to three orders of magnitude below the relation. This is particularly true for pulsars with both higher and lower weak spin-down rates than those among the Lyne et al. (2010) sample.
@@ 5.2 Implications for precision timing experiments
Millisecond pulsars are known to display emission state switching on short timescales (Mahajan et al., 2018; Miles et al., 2022; Nathan et al., 2023) and transient profile shape changes (Shannon et al., 2016; Lam et al., 2018; Goncharov et al., 2021; Jennings et al., 2024), with the latter effect requiring either profile-domain timing techniques to correct for them or careful removal of the corresponding signal in the timing residuals (see, e.g. Shannon et al. 2016). The propensity of spin-down rate variations among our pulsar sample and clear positive scaling in variability amplitude with suggests this phenomena should also be both present and detectable among the millisecond pulsar population. Accurately correcting for these effects is particularly important for precision timing experiments aiming to detect the stochastic gravitational-wave background.
@@
Assuming the rotational characteristics of millisecond pulsars follows the same scaling as the young pulsar population, then our results can be used to predict the expected spin-down rate variability among the pulsars currently monitored by pulsar timing arrays. The spin-down rates of the pulsars included in the upcoming third data release of the International Pulsar Timing Array (K. Liu, private communication; Zic et al. 2023; EPTA Collaboration et al. 2023a; Agazie et al. 2023b; Miles et al. 2025a) largely range between s-2 and s-2, alongside several noteworthy outliers such as PSR B193721 with its high spin-down rate of s-2. If the variable spin-down rate process is active among them, then from Figure 12 we would expect these pulsars to display changes in with amplitudes ranging between to s-2. This would be easily discernable using the Gaussian process regression framework that we employed.
@@ 5.3 Quasi-periodic variability
Among our broader sample there exists a sub-group of 45 pulsars that display striking quasi-periodic oscillations in their spin-down timeseries. These pulsars were predominantly identified through a combination of visual inspection of their timing residuals, spin-down timeseries and Lomb-Scargle periodograms. This group includes five that were previously identified by Lyne et al. (2010) and Kerr et al. (2016). Modulation periods () listed in Table 1 were recovered from Lomb-Scargle periodograms generated from their spin-down timeseries. We failed to recover the sub-annual periodicities of 120 d and 123 d in PSRs J16464346 and J18251446 that were found by Kerr et al. (2016). This may be due to a combination of our extended timing baselines and differences in methodology. Ten of the pulsars in Table 1 display both long and short modulation timescales that may be harmonically related. The ‘long’ modulation periods in PSRs J11365525, J12436423, J16494653, J17054306 and J18301059 are either close to or exactly twice the duration of the corresponding ‘short’ periods in these pulsars, and hence may be second harmonics. In comparison the dual modulation timescales in PSRs J10566258 J12246407, J13266700, J16005751 and J17223712 display approximate fractional relations of 5/2, 7/4, 9/5, 3/2 and 7/5 respectively.
@@
The lack of consistency between our much larger sample of pulsars with quasi-period variations and these predicted relations suggests that simple models of elastic strain or superconductivity driven free precession are not the dominant drivers of this behaviour. Jones (2012) noted that the relationships with required all pulsars display similar levels of strain in their crusts, while superconductivity necessitated that the entire star undergo precession. Relaxing these assumptions would resolve these two issues, but would be difficult to falsify. Another complicating factor is the presence of glitches among many of these pulsars, which can arise from interactions between the internal superfluid and the crust. Pinning of superfluid vortices within the crust is expected to dampen free-precession on relatively short timescales (Jones & Andersson, 2001; Link & Epstein, 2001; Jones et al., 2017). Non-radial oscillation modes of varying velocities and amplitudes have been suggested as an alternate means of driving changes in pulsar emission and spin-down, though without a clear mechanism (Rosen et al., 2011). Hall waves launched by transient events both in the superfluid core and crust of neutron stars have been recently proposed as a means of generating long-term quasi-periodic oscillations in that affect the inferred pulsar braking index (Bransgrove et al., 2024). The periods of the waves investigated are significantly longer (1-100 kyr) than the quasi-periodic processes we observe, but shorter period waves could be launched by starquakes acting as glitch triggers (Bransgrove et al., 2020).
@@ 5.4 Transient events and interactions with planetesimals
Out of the 238 pulsars in Figure 5 with variable spin-down rates, we were able to visually identify 68 individual transient spin-down events in 26 pulsars. The absolute size and duration of these events along with the inferred asteroid masses are listed in Table 4. We caution that this sample is incomplete to small transient changes in , which are inherently difficult to distinguish from other stochastic variations and sinusoidal changes in arising from residual positional offsets. Comparing the putative asteroid masses, the smallest resulted in the second event in PSR J18330827 with g. The third (MJD 54600-55400) event in PSR J16025100 is the largest with a mass of g. This is however smaller than the g mass inferred for the large 2005 spin-down/profile event detected in PSR J07384042 (Brook et al., 2014; Lower et al., 2023). Under the simplifying assumption that these objects were approximately spherical and C-type (carbonaceous) asteroids with a density of g cm-3 (Carry, 2012), the computed diameters range between to cm and cm for the 2005 event in PSR J07384042. Objects within this size range are exceedingly common among main-belt asteroids in the Solar System (Durda et al., 1998). If these transient spin-down events are indeed the result of pulsar-asteroid interactions, then their prevalence implies that debris discs and asteroids orbiting pulsars should be relatively common. Previous searches for planetesimals orbiting pulsars have largely been unsuccessful, though the lowest accessible mass ranges of these searches have been predominantly limited by the achieved timing precision. Surveys of 100’s of slow pulsars placed stringent limits on the presence of orbiting bodies larger than the mass of Mercury (Kerr et al., 2015; Niţu et al., 2022), while analyses of millisecond pulsar timing arrays have pushed these limits to between - Lunar masses (Behrens et al., 2020; Niţu et al., 2024). These limits are several orders of magnitude higher than the putative asteroid masses listed in Table 4.
@@
Gravitational scattering of ISOs on hyperbolic encounters can produce cuspy signals in pulsar timing residuals, where the sharpness and sign of the cusp depends on the viewing geometry and periastron distance, and the amplitude on the ISO mass (Jennings et al., 2020a). Close encounters of around 1 AU can produce signals that can appear indistinguishable from a glitch given a low observation cadence. More distant interactions would result in a more rounded peak, strikingly similar to those associated with the timing events in several pulsars displayed in Figure 2. However, given the linear scaling between the timing perturbation and ISO mass (Equation 16 in Jennings et al. 2020a), we can infer that interactions ISOs with masses equivalent to that of gas giant planets (i.e M⊙) would be required to reproduce the peaks of between 10’s-100’s of milliseconds seen among our pulsar sample.. It is also unclear how such distant encounters would alter the radio emission mechanism of the pulsar.
@@ 6 Conclusions
Alongside this, we showed that the various expected scaling relations from models of free-precession are unable to fully account for the large spread in modulation periods seen among a sub-group of 45 pulsars that display extremely (quasi-)periodic spin-down rate (and occasionally, profile shape) variations. This lends weight towards a magnetospheric state-switching origin for the quasi-periodic effects. We also re-examined interactions between pulsars and infalling planetesimals as a means of producing transient changes in profile and spin-down rate among 26 pulsars. Direct collisions and hyperbolic flybys of interstellar objects are ruled out, with the rate of transient spin-down events among these pulsars being four orders of magnitude higher than the theoretical impact rate, whereas repeat flybys of Jupiter-mass objects are required to explain the multiple large transient events seen in individual pulsars. Disruption of infalling asteroids from debris disks orbiting these pulsars remains a valid explanation. Assuming C-type composition asteroids, the masses and radii inferred from the amplitude and duration of these events are consistent with that are prevalent in our Solar System. Future direct detections of thermal emission from asteroid belts and debris disks around these pulsars via deep infrared or sub-mm observations would lend substantial weight towards this hypothesis. At present, a purely magnetospheric origin cannot be ruled out.
Alongside this, we showed that free precession, a popular mechanism for producing highly quasi-periodic changes in profile shape and spin-down, is unable to fully account for the spread in quasi-periods given other observed pulsar properties. We also rule out interactions between pulsars and interstellar objects as a means of generating transient profile/spin-down events. Though interactions with orbiting asteroids and magnetospheric fluctuations both remain viable mechanisms for explaining these effects.
An apparent deficit in the number of pulsars displaying profile shape changes, despite showing clear variations, is largely due to a combination of relatively low per-epoch signal-to-noise ratios, short observations resulting in stochastic jitter-dominated shape changes and smearing of profile components by scatter-broadening in the interstellar medium. Similar challenges to identifying profile-shape changing pulsars have been faced by other large-scale timing programmes, including the Thousand Pulsar Array on MeerKAT where few new shape changing pulsars have been identified to date despite the substantially higher telescope gain over Murriyang (Basu et al., 2024). Increasing this population could be achieved through a re-optimisation of existing programmes and those proposed to take place using the SKA. Increasing the duration of each pulsar observation would reduce the impact of pulse-to-pulse jitter, and make such programmes sensitive to more subtle shape changes.
@@ Appendix A Spin-down model parameters and inferred magnetospheric properties
Truncated tables listing the best fit model parameters for the timing residual Gaussian process regression fits, pulsar rotational and derived magnetospheric parameters, and transient spin-down event/asteroid properties. The full tables in both .pdf and machine-readable .csv format can be found in the Supplementary Materials.

### v3

@@ 4.2 New pulsars with both profile and spin-down variations
The radio profile of this pulsar is has two peaks that we found to exhibit substantial variability. The leading component shows both fast and slow changes, while the trailing component only varies slowly with time. In contrast, the spin-down rate varies smoothly over time, with a potential quasi-period of 1648 d. Our correlation map suggests that the profile variations lag the changing spin-down rate by 500-1000 d.
@@ 5.1 Scaling relations across the population
While a direct causal relationship between the observed spin-down variability in pulsars and external or internal physical processes is yet to be established, we can explore how this behaviour scales phenomenologically with various pulsar properties. In Figure 12 we present the measured min/max difference in spin-down rate against the corresponding ‘weak’ (minimum) spin-down rate value for the 238 variable pulsars in our sample, alongside values from several previous works. Lyne et al. (2010) noted that their sample of 17 correlated profile/spin-down state switching pulsars appeared to follow a simple power-law relation, where . A somewhat shallower power-law relation of was recovered using extended timing of the same pulsar sample by Shaw et al. (2022). While there is a substantial number of pulsars in Figure 12 that reside about the original Lyne et al. (2010) relation (shown in black), about half of our larger sample are situated between one to three orders of magnitude below the relation. This is particularly true for pulsars with both higher and lower weak spin-down rates than those among the Lyne et al. (2010) sample.
@@ 5.2 Implications for precision timing experiments
Millisecond pulsars are known to display emission state switching on short timescales (Mahajan et al., 2018; Miles et al., 2022; Nathan et al., 2023) and transient profile shape changes (Shannon et al., 2016; Lam et al., 2018; Goncharov et al., 2021; Jennings et al., 2024), with the latter effect requiring either profile-domain timing techniques to correct for them or careful removal of the corresponding signal in the timing residuals (see, e.g. Shannon et al. 2016). The propensity of spin-down rate variations among our pulsar sample and clear positive scaling in variability amplitude with suggests that this phenomenon may also be both present and detectable among the millisecond pulsar population. Accurately correcting for these effects is particularly important for precision timing experiments aiming to detect the stochastic gravitational-wave background.
@@
Assuming the rotational characteristics of millisecond pulsars follow the same scaling as the young pulsar population, our results can be used to predict the expected spin-down rate variability among the pulsars currently monitored by pulsar timing arrays. The spin-down rates of the pulsars included in the upcoming third data release of the International Pulsar Timing Array (K. Liu, private communication; Zic et al. 2023; EPTA Collaboration et al. 2023a; Agazie et al. 2023b; Miles et al. 2025a) largely range between s-2 and s-2, alongside several noteworthy outliers such as PSR B193721 with its high spin-down rate of s-2. If the variable spin-down rate process is active among them, then from Figure 12 we would expect these pulsars to display changes in with amplitudes ranging between to s-2. This would be easily discernable using the Gaussian process regression framework that we employed.
@@ 5.3 Quasi-periodic variability
Among our broader sample there exists a sub-group of 45 pulsars that display striking quasi-periodic oscillations in their spin-down timeseries. These pulsars were predominantly identified through a combination of visual inspection of their timing residuals, spin-down timeseries and Lomb-Scargle periodograms. This group includes five that were previously identified by Lyne et al. (2010) and Kerr et al. (2016). Modulation periods () listed in Table 1 were recovered from Lomb-Scargle periodograms generated from their spin-down timeseries. We failed to recover the sub-annual periodicities of 120 d and 123 d in PSRs J16464346 and J18251446 that were found by Kerr et al. (2016). This may be due to a combination of our extended timing baselines and differences in methodology. Ten of the pulsars in Table 1 display both long and short modulation timescales that may be harmonically related. The ‘long’ modulation periods in PSRs J11365525, J12436423, J16494653, J17054306 and J18301059 are either close to or exactly twice the duration of the corresponding ‘short’ periods in these pulsars. Hence, the short periods may be second harmonics. In comparison the dual modulation timescales in PSRs J10566258 J12246407, J13266700, J16005751 and J17223712 display approximate fractional relations of 5/2, 7/4, 9/5, 3/2 and 7/5 respectively.
@@
The lack of consistency between our much larger sample of pulsars with quasi-period variations and these predicted relations suggests that simple models of elastic strain or superconductivity driven free precession are not the dominant drivers of this behaviour. Jones (2012) noted that the relationships with required all pulsars display similar levels of strain in their crusts, while superconductivity necessitated that the entire star undergo precession. Relaxing these assumptions would resolve these two issues, but would be difficult to falsify. Another complicating factor is the presence of glitches among many of these pulsars, which can arise from interactions between the internal superfluid and the crust. Pinning of superfluid vortices within the crust is expected to dampen free-precession on relatively short timescales (Jones & Andersson, 2001; Link & Epstein, 2001; Jones et al., 2017). Non-radial oscillation modes of varying velocities and amplitudes have been suggested as an alternate means of driving changes in pulsar emission and spin-down, though without a clear mechanism (Rosen et al., 2011). Hall waves launched by transient events both in the superfluid core and crust of neutron stars have been recently proposed as a means of generating long-term quasi-periodic oscillations in that affect the inferred pulsar braking index (Bransgrove et al., 2025). The periods of the waves investigated are significantly longer (1-100 kyr) than the quasi-periodic processes we observe, but shorter period waves could be launched by starquakes acting as glitch triggers (Bransgrove et al., 2020).
@@ 5.4 Transient events and interactions with planetesimals
Out of the 238 pulsars in Figure 5 with variable spin-down rates, we were able to visually identify 68 individual transient spin-down events in 26 pulsars. The absolute size and duration of these events along with the inferred asteroid masses are listed in Table 10. We caution that this sample is incomplete to small transient changes in , which are inherently difficult to distinguish from other stochastic variations and sinusoidal changes in arising from residual positional offsets. Comparing the putative asteroid masses, the smallest resulted in the second event in PSR J18330827 with g. The third (MJD 54600-55400) event in PSR J16025100 is the largest with a mass of g. This is however smaller than the g mass inferred for the large 2005 spin-down/profile event detected in PSR J07384042 (Brook et al., 2014; Lower et al., 2023). Under the simplifying assumption that these objects were approximately spherical and C-type (carbonaceous) asteroids with a density of g cm-3 (Carry, 2012), the computed diameters range between to cm and cm for the 2005 event in PSR J07384042. Objects within this size range are exceedingly common among main-belt asteroids in the Solar System (Durda et al., 1998). If these transient spin-down events are indeed the result of pulsar-asteroid interactions, then their prevalence implies that debris discs and asteroids orbiting pulsars should be relatively common. Previous searches for planetesimals orbiting pulsars have largely been unsuccessful, though the lowest accessible mass ranges of these searches have been predominantly limited by the achieved timing precision. Surveys of hundreds of slow pulsars placed stringent limits on the presence of orbiting bodies larger than the mass of Mercury (Kerr et al., 2015; Niţu et al., 2022), while analyses of millisecond pulsar timing arrays have pushed these limits to between - Lunar masses (Behrens et al., 2020; Niţu et al., 2024). These limits are several orders of magnitude higher than the putative asteroid masses listed in Table 10.
@@
Gravitational scattering of ISOs on hyperbolic encounters can produce cuspy signals in pulsar timing residuals, where the sharpness and sign of the cusp depends on the viewing geometry and periastron distance, and the amplitude on the ISO mass (Jennings et al., 2020a). Close encounters of around 1 AU can produce signals that can appear indistinguishable from a glitch given a low observation cadence. More distant interactions would result in a more rounded peak, strikingly similar to those associated with the timing events in several pulsars displayed in Figure 2. However, given the linear scaling between the timing perturbation and ISO mass (Equation 16 in Jennings et al. 2020a), we can infer that interactions ISOs with masses equivalent to that of gas giant planets (i.e M⊙) would be required to reproduce the peaks of between 10’s-100’s of milliseconds seen among our pulsar sample. It is also unclear how such distant encounters would alter the radio emission mechanism of the pulsar.
@@ 6 Conclusions
Alongside this, we showed that the various expected scaling relations from models of free-precession, a popular mechanism for producing highly quasi-periodic changes in profile shape and spin-down, are unable to fully account for the large spread in modulation periods seen among a sub-group of 45 pulsars that display extremely (quasi-)periodic spin-down rate (and occasionally, profile shape) variations. This lends weight towards a magnetospheric state-switching origin for the quasi-periodic effects. We also re-examined interactions between pulsars and infalling planetesimals as a means of producing transient changes in profile and spin-down rate among 26 pulsars. Direct collisions and hyperbolic flybys of interstellar objects are ruled out, with the rate of transient spin-down events among these pulsars being four orders of magnitude higher than the theoretical impact rate, whereas repeat flybys of Jupiter-mass objects are required to explain the multiple large transient events seen in individual pulsars. Disruption of infalling asteroids from debris disks orbiting these pulsars remains a valid explanation. Assuming C-type composition asteroids, the masses and radii inferred from the amplitude and duration of these events are consistent with planetesimals that are prevalent in our Solar System. Future direct detections of thermal emission from asteroid belts and debris disks around these pulsars via deep infrared or sub-mm observations would lend substantial weight towards this hypothesis. A purely magnetospheric origin for these behaviours cannot be ruled out at present.
An apparent deficit in the number of pulsars displaying profile shape changes, despite showing clear variations, is largely due to a combination of relatively low per-epoch signal-to-noise ratios, short observations resulting in stochastic jitter-dominated shape changes and smearing of profile components by scatter-broadening in the interstellar medium. Similar challenges to identifying profile-shape changing pulsars have been faced by other large-scale timing programmes. This includes the Thousand Pulsar Array on MeerKAT, where few new shape-changing pulsars have been identified to date despite the substantially higher telescope gain over Murriyang (Basu et al., 2024). Increasing this population could be achieved through a re-optimisation of existing programmes and those proposed to take place using the SKA. Lengthening the duration of each pulsar observation would reduce the impact of pulse-to-pulse jitter, and make such programmes sensitive to more subtle shape changes.
@@ Appendix A Spin-down model parameters and inferred magnetospheric properties
Tables listing the best fit model parameters for the timing residual Gaussian process regression fits, pulsar rotational and derived magnetospheric parameters, and transient spin-down event/asteroid properties.